# Performance and Storage Analysis of CRYSTALS-Kyber (ML-KEM) as a Post-Quantum Replacement for RSA and ECC

### v2

@@ 5.1.1 Key Generation
Kyber demonstrates superior efficiency in key generation. On both x86_64 and ARM64 architectures, Kyber requires the fewest cycles ( 7.4M and  7.3M, respectively), making it approximately 2.7 to 3 times faster than SECP384R1 ( 21.3M and  19.5M cycles). The performance gap with RSA is staggering. On the x86_64 platform, Kyber is over 20,500 times faster than RSA, and on ARM64, it remains over 3,400 times faster. This vast difference is due to the underlying mathematics of the algorithms. RSA key generation relies on the computationally intensive and time-consuming process of finding large prime numbers, which scales poorly. In contrast, Kyber’s lattice-based arithmetic allows for much more efficient key creation. Although the ARM64 architecture significantly reduces RSA’s key generation time by a factor of six compared to x86_64, it remains orders of magnitude slower than both Kyber and ECC.
@@ 5.1.2 Outgoing Shared Secret Derivation
The data reveals that Kyber is the most efficient algorithm for this operation on both tested platforms. On x86_64, Kyber’s cost of approximately 210,800 cycles is roughly 25 times faster than RSA ( 5.2 million cycles) and 72 times faster than SECP384R1 ( 15.2 million cycles). On the ARM64 architecture, Kyber ( 288,200 cycles) maintains its lead, proving to be about 13 times faster than RSA ( 3.7 million cycles) and 47 times faster than SECP384R1 ( 13.6 million cycles).
It is important to note that for RSA, this “public-key” operation is significantly less costly than the “private-key” operation shown in the “Incoming Derivation” chart, which is an expected characteristic of the algorithm. Nevertheless, Kyber still surpasses it by a substantial margin. For SECP384R1, the computational workload is identical for both the incoming and outgoing phases, which explains its consistently high cycle count. Kyber’s efficiency, requiring fewer than 300,000 cycles, facilitates near-instantaneous symmetric key establishment, drastically reducing the latency of cryptographic handshakes compared to both RSA and ECC.
@@ 5.1.3 Incoming Shared Secret Derivation
On the x86_64 architecture, Kyber requires only about 236,700 cycles. This is approximately 64 times more efficient than SECP384R1, which takes 15.2 million cycles, and over 3,200 times more efficient than RSA, which needs 764 million cycles for the same task. A similar trend is observed on the ARM64 platform, where Kyber’s 329,900 cycles outperform SECP384R1 (13.6 million cycles) by a factor of 41 and RSA (529.6 million cycles) by a factor of over 1,600.
It is important to note that for
SECP384R1, the derivation procedure is identical for both the outgoing and incoming phases, which explains its consistently high computational cost compared to Kyber in both scenarios. This gap stems from Kyber’s foundation in module lattice-based arithmetic, which avoids the expensive elliptic curve point multiplication used by SECP384R1 and the even more costly modular exponentiation of RSA.
@@ 5.2 Storage benchmarks
The added overhead refers to the data that an algorithm adds to the existing data (plaintext length) to encrypt it, and in some cases, to include the key transport, for example, with Kyber, is: plaintext message length + Kyber’s overhead + Rust’s serialization bytes. The storage overhead analysis reveals a critical trade-off between quantum resistance and bandwidth efficiency. While the Poly1305 authentication nonce contributes a constant 30-byte overhead across all schemes, the key transport mechanism diverges significantly: SECP384R1 (ECC) requires zero key transport overhead due to its elliptic curve key agreement protocol, resulting in a minimal total added length of 30 bytes. In contrast, Kyber and RSA introduce substantial key transport payloads (1,091 bytes and 963 bytes, respectively), resulting in total overheads of 1,121 bytes for Kyber and 993 bytes for RSA. This establishes ECC as the most bandwidth-efficient scheme—advantageous in constrained environments, such as IoT—but its quantum vulnerability remains an existential risk. Kyber’s $13.5\%$ higher transport overhead than RSA is justified by its quantum resistance and superior computational efficiency (see Section 5.1), positioning it as the optimal choice for future-proof systems where bandwidth permits.
@@ Conclusion
These highly positive results, obtained using built-in CPU features like AVX2 and AES-NI, strongly suggest a path for future innovation: the design and development of specialized hardware accelerators for lattice-based cryptography. The creation of such hardware could unlock further performance gains, reducing latency and power consumption to solidify the viability of post-quantum cryptography in even the most resource-constrained environments, such as IoT devices.

### v4

@@ 5.1.1 Key Generation
Kyber shows efficient key generation. On both x86_64 and ARM64 architectures, Kyber requires the fewest cycles ( 7.4M and  7.3M, respectively), making it approximately 2.7 to 3 times faster than SECP384R1 ( 21.3M and  19.5M cycles). In comparison with RSA, the difference is substantial: on the x86_64 platform Kyber uses roughly 20,500 times fewer cycles, and on ARM64 about 3,400 times fewer. These factors are consistent with the underlying complexity of each algorithm. RSA key generation involves searching for large prime numbers, which is computationally intensive and scales poorly, whereas Kyber’s lattice-based arithmetic admits more efficient implementations. Although the ARM64 architecture significantly reduces RSA’s key generation time by a factor of six compared to x86_64, it is still much slower than both Kyber and ECC.
@@ 5.1.2 Outgoing Shared Secret Derivation
Kyber is the most efficient algorithm for this operation on both tested platforms. On x86_64, Kyber’s cost of approximately 210,800 cycles is roughly 25 times lower than RSA ( 5.2 million cycles) and about 72 times lower than SECP384R1 ( 15.2 million cycles). On the ARM64 architecture, Kyber ( 288,200 cycles) maintains its lead, remaining roughly 13 times more efficient than RSA ( 3.7 million cycles) and 47 times more efficient than SECP384R1 ( 13.6 million cycles).
It is important to note that for RSA, this “public-key” operation is significantly less costly than the “private-key” operation shown in the “Incoming Derivation” chart, which is an expected characteristic of the algorithm. Nevertheless, Kyber’s performance remains superior. For SECP384R1, the computational workload is identical for both the incoming and outgoing phases, which explains its consistently high cycle count. Kyber requires fewer than 300,000 cycles for this step, which reduces the latency of cryptographic handshakes relative to both RSA and ECC.
@@ 5.1.3 Incoming Shared Secret Derivation
On the x86_64 architecture, Kyber requires about 236,700 cycles for incoming shared secret derivation, whereas SECP384R1 needs 15.2 million cycles and RSA around 764 million cycles. The resulting factors (roughly 64× and 3,200×, respectively) arise from the different mathematical operations underlying each algorithm. A similar trend is observed on the ARM64 platform, where Kyber’s 329,900 cycles outpace SECP384R1 (13.6 million cycles) by about 41× and RSA (529.6 million cycles) by about 1,600×.
For SECP384R1, the derivation procedure is identical for both the outgoing and incoming phases, which explains its consistently high computational cost relative to Kyber in both scenarios. These differences follow from Kyber’s foundation in module lattice-based arithmetic, which avoids the expensive elliptic curve point multiplication used by SECP384R1 and the even more costly modular exponentiation of RSA.
@@ 5.2 Storage benchmarks
The added overhead refers to the data that an algorithm adds to the existing data (plaintext length) to encrypt it, and in some cases, to include the key transport, for example, with Kyber, is: plaintext message length + Kyber’s overhead + Rust’s serialization bytes. The storage overhead analysis highlights a trade-off between post-quantum security and bandwidth. The Poly1305 authentication nonce adds a constant 30-byte overhead across all schemes, but the key transport payload varies: SECP384R1 uses elliptic curve key agreement and thus incurs no extra key transport overhead, yielding a total added length of 30 bytes. Kyber and RSA require additional key transport data (1,091 bytes and 963 bytes, respectively), leading to total overheads of 1,121 bytes for Kyber and 993 bytes for RSA. ECC’s minimal overhead makes it the most bandwidth-efficient option—a consideration for constrained environments such as IoT—but its lack of quantum resistance limits its longevity. Kyber’s roughly $13.5\%$ larger transport payload than RSA reflects the size of its lattice-based ciphertexts; given its computational advantages and quantum security, this overhead is likely acceptable for many applications.
@@ Conclusion
These findings suggest that, with existing CPU support for instructions like AVX2 and AES-NI, lattice-based cryptography is practical today. Future development of specialized hardware accelerators for lattice operations could further reduce latency and power consumption, expanding the range of applications—including resource-constrained environments such as IoT devices—where post-quantum cryptography may be deployed.